OpenStudy (anonymous):

Studying for Calc quiz,please help me figure out this limit. Lim t approaches infinity squareroot t + t^2 / 3t-t^2. I put real equation in commentss

OpenStudy (anonymous):

\[ \lim_{t \rightarrow \infty} \frac{ \sqrt{t}+t^2 }{ 3t-t^2 }\]